OpenStudy (ihelpyoulearn):

Solve the system of equations algebraically. Show all your steps if you can please that way I can learn. Y=x^2+2x Y=3x+20 I appreciate any help 😊 I tried everything I could think of but I just can't get it right! I will medal anyone who gives effort to help me. Thank You!!!

OpenStudy (gniiij):

i can help

OpenStudy (ihelpyoulearn):

Really? That would be great since I can't seem to be doing it right

OpenStudy (gniiij):

so what you would want to do it substitute the value of y from 1 equation into the value of y in the second equation

OpenStudy (gniiij):

for example: 3x+20=x^2+2x

OpenStudy (ihelpyoulearn):

Yes, I've got that 😊

OpenStudy (gniiij):

20=x^2+2x-3x

OpenStudy (gniiij):

20=x^2-x

OpenStudy (gniiij):

0=x^2-x-20

OpenStudy (gniiij):

then you want to factor x^2-x-20

OpenStudy (gniiij):

so can you factor that for me?

OpenStudy (gniiij):

you there?

OpenStudy (ihelpyoulearn):

@GNoiij Every time I try to factor this part I get x=-4 or x=5?

OpenStudy (ihelpyoulearn):

Oops my bad. @GNiiij